Check whether the sequence, {Sn } WHERE
,Sn =1/1!+1/3!+1/5!+...+1/(2n-1)!


Is convergent or not.
1
Expert's answer
2020-04-16T15:15:49-0400

Use the Ratio Test


"\lim\limits_{n\to\infin}\big|{a _{n+1}\over a_n}\big|=\lim\limits_{n\to\infin}\bigg|{\dfrac{1}{(2(n+1)-1)!}\over \dfrac{1}{(2n-1)!}}\bigg|="

"=\lim\limits_{n\to\infin}\big|{1\over 2n(2n+1)}\big|=0"

The sequence is convergent by the Ratio Test.